• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Wendy

    Log entry time 02:44:40 on November 2,2001

    Entry number 75602

    This entry is a followup to: 75546


    keyword=Target Issue

    Mattias and I looked at different histograms in an attempt to figure out what went wrong with the target earlier. The detector histograms for the runs 1404 (before), 1410 (during), and 1420 (after) all seem to be fine. The figure is of y target as seen by the electron spectrometer. The shaded plots are 1404 and 1420 and the outlined plots are 1410 and 1411. Scattering from the Be windows is increased for 1410 and 1411, while the scattering from the water is less for these two runs than it was for 1404 and 1420. (The plots have been normalized to charge.)

    FIGURE 1